# Dirac Fermions and Topological Phases in Magnetic Topological Insulator Films

### v2

@@ III.3 A glance in proof of half-quantization
where $A^{M}$ refers to the non-Abelian Berry connection (convention follows that $\mathcal{A}=i\braket{u}{\mathrm{d}}{u}$) formed by the metallic bands crossed by the Fermi surface with parity or time-reversal symmetry, while $A^{L}$ refers to connection of bands with lower energy, on the boundary formed by $k_{F}$. Essentially, the last two terms are phase integrals around one mutual boundary with opposite orientations, which will contribute an integer value[71, 72, 73] $2\pi C$. For the first term, requiring the 2D parity (i.e., mirror) symmetry at the Fermi surface leads to a local unitary transformation $U^{M}_{k}$ relating states at parity-symmetric points, which leads to
@@ IV.1 Magnetism polarized along $z$ direction
which acts on spin $z$. For several schematic examples with different Zeeman configurations, see Fig. 8. Its equivalent action by projection $\braket{\Phi^{n}_{m}}{V_{Z}}{\Phi^{n^{\prime}}_{m^{\prime}}}$ ($m,m^{\prime}=1,2,3,4$; $n,n^{\prime}=1,\cdots,L_{z}$) reads
@@
where $n,n^{\prime}=1,\cdots,L_{z}$. Notice that $I_{S/A}$ is non-vanishing only when the symmetric/antisymmetric component of $V_{z}$ is non-zero. Our formula then illustrates that the Zeeman field in a TI film is brought into two classes by the discrete parity or mirror symmetry, with $S$($A$) labelling the part respects (disrespects) this symmetry. Bring the transformed Zeeman term into multi-Dirac fermions representation, and we obtain
@@ IV.1.2 Weak Zeeman field
$\begin{cases}I_{S}(\bm{k})&=\braket{\Phi_{1}(\bm{k})}{V_{Z}}{\Phi_{1}(\bm{k})}\\ iI_{A}(\bm{k})&=\braket{\Phi_{1}(\bm{k})}{V_{Z}}{\Phi_{3}(\bm{k})}\end{cases},$ (72)
@@ IV.2 Other fields
Zeeman field $Z_{z}(l_{z})\sigma_{z}\tau_{0}$ $\left(\mathbf{I}^{z}_{S}(\bm{k})\tau_{0}-\mathbf{I}^{z}_{A}(\bm{k})\tau_{y}\right)\sigma_{z}$ $F_{S+},\leavevmode\nobreak\ F_{A+}$
$Z_{x}(l_{z})\sigma_{x}\tau_{0}$ $\left(\mathbf{I}^{x}_{S}(\bm{k})\tau_{x}-\mathbf{I}^{x}_{A}(\bm{k})\tau_{z}\right)\sigma_{x}$ $F_{S-},\leavevmode\nobreak\ F_{A-}$
$Z_{y}(l_{z})\sigma_{y}\tau_{0}$ $\left(\mathbf{I}^{y}_{S}(\bm{k})\tau_{x}-\mathbf{I}^{y}_{A}(\bm{k})\tau_{z}\right)\sigma_{y}$ $F_{S-},\leavevmode\nobreak\ F_{A-}$
Gate-voltage $G(l_{z})\sigma_{0}\tau_{0}$ $\left(\mathbf{G}_{S}(\bm{k})\tau_{0}-\mathbf{G}_{A}(\bm{k})\tau_{y}\right)\sigma_{0}$ $F_{S+},\leavevmode\nobreak\ F_{A+}$
Oribital field $O_{y}(l_{z})\sigma_{0}\tau_{y}$ $\left(\mathbf{O}^{y}_{A}(\bm{k})\tau_{0}-\mathbf{O}^{y}_{S}(\bm{k})\tau_{y}\right)\sigma_{z}$ $F_{S+},\leavevmode\nobreak\ F_{A+}$
$O_{x}(l_{z})\sigma_{0}\tau_{x}$ $\left(\mathbf{O}^{x}_{A}(\bm{k})\tau_{z}-\mathbf{O}^{x}_{S}(\bm{k})\tau_{x}\right)\sigma_{0}$ $-F_{S-},\leavevmode\nobreak\ -F_{A-}$
$O_{z}(l_{z})\sigma_{0}\tau_{z}$ $\left(\mathbf{O}^{z}_{A}(\bm{k})\tau_{x}-\mathbf{O}^{z}_{S}(\bm{k})\tau_{z}\right)\sigma_{z}$ $F_{S-},\leavevmode\nobreak\ -F_{A-}$
For a given field $V(\bm{k},l_{z})\sigma_{\mu}\tau_{\nu}$, the transformation follows similarly by organizing the projected elements $\sum_{l_{z}}V(\bm{k},l_{z})\braket{\Phi^{n}_{m}(\bm{k},l_{z})}{\sigma_{\mu}\tau_{\nu}}{\Phi^{n^{\prime}}_{m^{\prime}}(\bm{k},l_{z})}$ ($m,m^{\prime}=1,2,3,4$; $n,n^{\prime}=1,\cdots,L_{z}$) aligned with the sequence of the basis. The form of field after transformation will always be two $L_{z}\times L_{z}$ matrix fields differing by $z$-parity symmetry labels, with $S$ counting for symmetric distribution and $A$ for the opposite. Each matrix field will also be attached with a new $4\times 4$ Dirac matrix.
To express matrix quantities $\mathbf{I},\mathbf{G},\mathbf{O}$ in Table 1, we introduce the momentum-dependent matrix-form acting functional $\mathcal{F}_{k}$ over $V$ field that generates projected matrix component like
@@
different by symmetry requirement and an inner sign. In the table the symmetry labels between the transformed fields and the summation kernels are in one-to-one correspondence.
The table can be longer once one considers more kinds of Dirac matrices. This procedure above is general, powerful while easy to understand. Despite the easiness of the transformation, the non-trivial difficult part is to endow physical meaning to the attached fields, both before transformation and after. For instance, the spin-orbital coupling remains its meaning after the transformation, while being block-diagonal in the Dirac fermion representation; the $z$-Zeeman field, as discussed above, is transformed into two matrix form Higgs fields, which stand as the effective mass generators.
@@ V.1 Half quantum mirror Hall effect: a non-magnetic film with mirror symmetry
where $\bar{v}_{i}^{mn,\chi}=\braket{n^{\chi}}{\partial_{k_{i}}H^{\chi}}{m^{\chi}}$ is the expectation value of the mirror velocity operator evaluated over eigenstates of the mirror-projected Hamiltonian. Clearly, this is just the usual Kubo formula[62] evaluated over the projected Hamiltonian $H_{\chi}$, and thanks to the imposed mirror symmetry, two parts with mirror label $\chi=\pm$ do not communicate with each other and are totally decoupled.
@@
To show the nature of the half quantum mirror Hall effect, we calculate the Hall conductivity of $H_{\chi}$ obtained from the mirror-projected TI film Hamiltonian, and of the split Dirac fermion $H_{\text{surf},\chi}$. The results are shown in Fig. 11, where the half-quantized transverse conductivity nature is shown for each $\chi$ part with inverse signs, indicating quantum spin Hall like physics[12, 82, 43, 44, 45, 83, 84, 85, 86], while the topological origin of the half-quantized mirror Hall conductivity is bound with the metallic gapless Dirac fermions[33]. Their massless low-energy parts distribute mirror-(anti-)symmetrically at both top and bottom surfaces of the TI film as a result from the bulk-boundary correspondence of 3D strong topological insulator[17], corresponding to states with mass ±$\pm\Theta(-m_{0}(\bm{k}))m_{0}(\bm{k})$ term at $m_{0}(\bm{k})>0$. Here, the symmetry statement is traced back to our basis, which is chosen to distribute along $z$ either mirror symmetrically or anti-symmetrically (see Appendix B). As a complete band, the surface Dirac cone does not end at a finite wavevector, but gradually emerges into the bulk with a regulated non-zero mass term represented by $\Theta(-m_{0}(\bm{k}))m_{0}(\bm{k})$ at $m_{0}(\bm{k})<0$, and it is this non-vanishing high-energy part that ultimately gives rise to the half-quantized Hall conductivity, as discussed in Section III, which finally reads by Eq. (40) as $\sigma_{H}^{\chi}=-\chi e^{2}/2h$, when the Fermi surface satisfies that $m_{0}(k_{F})>0$.
@@
and equals to quantum unit $-e^{2}/h$ in the case. The quantity reveals that, though, by opposite Hall conductivity, the charge current by a transverse electrical field vanishes as $\sigma_{H}=\sum_{\chi}\sigma_{H}^{\chi}=0$, the ‘mirror’ current does not, similar to that in quantum spin Hall effect. Nevertheless, a better way of looking at the half quantum mirror Hall effect may start from treating it as an intrinsic ‘spin’ Hall effect in metal, while the effect shows quantization with its transverse ‘spin’ Hall conductivity that shares a topological origin deeply related to the parity anomaly, and replacing ‘spin’ with ‘mirror’ leads to the observation that in different mirror sectors, the mirror current and the charge current will be either parallel or anti-parallel with the same quantized magnitude. Such a way of narration also lies in the lineage of induced dissipationless mirror current and dissipative longitudinal current, as they are both generated by metallic gapless Dirac fermions. To detect the mirror current, non-local electrical transport signals[87, 88, 89] are needed, while to reveal the quantized nature, one needs to perform a series of measurements to fully separate the dissipationless and dissipative currents[33], by changing the sample width and noticing the scale invariance of the Hall conductance.
@@ V.2 Quantum anomalous Hall effect: Chern Insulators
Figure 13: Chern insulator case I: total layer number $L_{z}=19$ with uniform Zeeman field $V_{z}\equiv 0.1\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$. (a) Comparison of band structure from TI film model and effective four-band Hamiltonian. (b) Calculated $I_{S}(\bm{k})$ and $I_{A}(\bm{k})$. (c) Calculated Hall conductivity from TI film model and effective four-band Hamiltonian. (d) Hall conductivity for $\chi=\pm$.
Figure 14: Chern insulator case II: total layer number $L_{z}=19$ with symmetric Zeeman field $V_{z}(l_{z})=0.1\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$ at top and bottom $2$ layers. (a) Comparison of band structure from the lowest four bands of TI film model and effective four-band Hamiltonian. (b) Calculated $I_{S}(\bm{k})$ and $I_{A}(\bm{k})$. (c) Calculated Hall conductivity from TI film model and effective four-band Hamiltonian. (d) Hall conductivity for $\chi=\pm$.
Figure 15: Chern insulator case III: total layer number $L_{z}=19$ with top-$2$-layer Zeeman field $V_{z}^{t}=0.1\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$ and bottom-$2$-layer field $V_{z}^{b}=0.2\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$. (a) Comparison of band structure from the lowest four bands of TI film model and effective four-band Hamiltonian. (b) Calculated $I_{S}(\bm{k})$ and $I_{A}(\bm{k})$. (c) Calculated Hall conductivity from TI film model and effective four-band Hamiltonian. (d) Hall conductivity for $\chi=\pm$.

### v3

@@ III.3 A glance in proof of half-quantization
where $A^{M}$ refers to the non-Abelian Berry connection (convention follows that $\mathcal{A}=i\braket{u|\mathrm{d}|u}$) formed by the metallic bands crossed by the Fermi surface with parity or time-reversal symmetry, while $A^{L}$ refers to connection of bands with lower energy, on the boundary formed by $k_{F}$. Essentially, the last two terms are phase integrals around one mutual boundary with opposite orientations, which will contribute an integer value[71, 72, 73] $2\pi C$. For the first term, requiring the 2D parity (i.e., mirror) symmetry at the Fermi surface leads to a local unitary transformation $U^{M}_{k}$ relating states at parity-symmetric points, which leads to
@@ IV.1 Magnetism polarized along $z$ direction
which acts on spin $z$. For several schematic examples with different Zeeman configurations, see Fig. 8. Its equivalent action by projection $\braket{\Phi^{n}_{m}|V_{Z}|\Phi^{n^{\prime}}_{m^{\prime}}}$ ($m,m^{\prime}=1,2,3,4$; $n,n^{\prime}=1,\cdots,L_{z}$) reads
@@
where $n,n^{\prime}=1,\cdots,L_{z}$. Notice that $I_{S/A}$ is non-vanishing only when the symmetric/antisymmetric component of $V_{z}$ is non-zero. Our formula then illustrates that the Zeeman field in a TI film is brought into two classes by the discrete parity or mirror symmetry, with SS(AA) labelling the part respects (disrespects) this symmetry. Bring the transformed Zeeman term into multi-Dirac fermions representation, and we obtain
@@ IV.1.2 Weak Zeeman field
$\begin{cases}I_{S}(\bm{k})&=\braket{\Phi_{1}(\bm{k})|V_{Z}|\Phi_{1}(\bm{k})}\\ iI_{A}(\bm{k})&=\braket{\Phi_{1}(\bm{k})|V_{Z}|\Phi_{3}(\bm{k})}\end{cases},$ (72)
@@ IV.2 Other fields
Zeeman field $Z_{z}(l_{z})\sigma_{z}\tau_{0}$ $\left(\mathbf{I}^{z}_{S}(\bm{k})\tau_{0}-\mathbf{I}^{z}_{A}(\bm{k})\tau_{y}\right)\sigma_{z}$ $F_{S+},\penalty 10000\ F_{A+}$
$\tilde{Z}_{z}(l_{z})\sigma_{z}\tau_{z}$ $\left(\tilde{\mathbf{I}}^{z}_{S}(\bm{k})\tau_{z}-\tilde{\mathbf{I}}^{z}_{A}(\bm{k})\tau_{x}\right)\sigma_{0}$ $F_{S-},F_{A-}$
$Z_{x}(l_{z})\sigma_{x}\tau_{z}$ $\left(\mathbf{I}^{x}_{S}(\bm{k})\tau_{y}-\mathbf{I}^{x}_{A}(\bm{k})\tau_{0}\right)\sigma_{y}$ $F_{S+},F_{A+}$
$\tilde{Z}_{x}(l_{z})\sigma_{x}\tau_{0}$ $\left(\tilde{\mathbf{I}}^{x}_{S}(\bm{k})\tau_{x}-\tilde{\mathbf{I}}^{x}_{A}(\bm{k})\tau_{z}\right)\sigma_{x}$ $F_{S-},\penalty 10000\ F_{A-}$
$Z_{y}(l_{z})\sigma_{y}\tau_{z}$ $\left(\mathbf{I}^{y}_{S}(\bm{k})\tau_{y}-\mathbf{I}^{y}_{A}(\bm{k})\tau_{0}\right)\sigma_{x}$ $F_{S+},F_{A+}$
$\tilde{Z}_{y}(l_{z})\sigma_{y}\tau_{0}$ $\left(\tilde{\mathbf{I}}^{y}_{S}(\bm{k})\tau_{x}-\tilde{\mathbf{I}}^{y}_{A}(\bm{k})\tau_{z}\right)\sigma_{y}$ $F_{S-},\penalty 10000\ F_{A-}$
Gate-voltage $G(l_{z})\sigma_{0}\tau_{0}$ $\left(\mathbf{G}_{S}(\bm{k})\tau_{0}-\mathbf{G}_{A}(\bm{k})\tau_{y}\right)\sigma_{0}$ $F_{S+},\penalty 10000\ F_{A+}$
Oribital field $O_{y}(l_{z})\sigma_{0}\tau_{y}$ $\left(\mathbf{O}^{y}_{A}(\bm{k})\tau_{0}-\mathbf{O}^{y}_{S}(\bm{k})\tau_{y}\right)\sigma_{z}$ $F_{S+},\penalty 10000\ F_{A+}$
$O_{x}(l_{z})\sigma_{0}\tau_{x}$ $\left(\mathbf{O}^{x}_{A}(\bm{k})\tau_{z}-\mathbf{O}^{x}_{S}(\bm{k})\tau_{x}\right)\sigma_{0}$ $F_{S-},\penalty 10000\ -F_{A-}$
$O_{z}(l_{z})\sigma_{0}\tau_{z}$ $\left(\mathbf{O}^{z}_{A}(\bm{k})\tau_{x}-\mathbf{O}^{z}_{S}(\bm{k})\tau_{z}\right)\sigma_{z}$ $F_{S-},\penalty 10000\ -F_{A-}$
Other
spin-orbital
field
$S_{xx}(l_{z})\sigma_{x}\tau_{x}$ $\left(\mathbf{S}_{S}^{xx}(\bm{k})\tau_{0}-\mathbf{S}_{A}^{xx}(\bm{k})\tau_{y}\right)\sigma_{x}$ $F_{S+},F_{A+}$
$S_{xy}(l_{z})\sigma_{x}\tau_{y}$ $\left(\mathbf{S}_{S}^{xy}(\bm{k})\tau_{z}-\mathbf{S}_{A}^{xy}(\bm{k})\tau_{x}\right)\sigma_{y}$ $F_{S-},F_{A-}$
$S_{yx}(l_{z})\sigma_{y}\tau_{x}$ $\left(\mathbf{S}_{S}^{yx}(\bm{k})\tau_{0}-\mathbf{S}_{A}^{yx}(\bm{k})\tau_{y}\right)\sigma_{y}$ $F_{S+},F_{A+}$
$S_{yy}(l_{z})\sigma_{y}\tau_{y}$ $\left(\mathbf{S}_{S}^{yy}(\bm{k})\tau_{z}-\mathbf{S}_{A}^{yy}(\bm{k})\tau_{x}\right)\sigma_{x}$ $F_{S-},F_{A-}$
$S_{zx}(l_{z})\sigma_{z}\tau_{x}$ $\left(\mathbf{S}^{zx}_{A}(\bm{k})\tau_{z}-\mathbf{S}^{zx}_{S}(\bm{k})\tau_{x}\right)\sigma_{z}$ $F_{S-},\penalty 10000\ -F_{A-}$
$S_{zy}(l_{z})\sigma_{z}\tau_{y}$ $\left(\mathbf{S}^{zy}_{A}(\bm{k})\tau_{0}-\mathbf{S}^{zy}_{S}(\bm{k})\tau_{y}\right)\sigma_{0}$ $F_{S+},\penalty 10000\ F_{A+}$
For a given field $V(\bm{k},l_{z})\sigma_{\mu}\tau_{\nu}$, the transformation follows similarly by organizing the projected elements $\sum_{l_{z}}V(\bm{k},l_{z})\braket{\Phi^{n}_{m}(\bm{k},l_{z})|\sigma_{\mu}\tau_{\nu}|\Phi^{n^{\prime}}_{m^{\prime}}(\bm{k},l_{z})}$ ($m,m^{\prime}=1,2,3,4$; $n,n^{\prime}=1,\cdots,L_{z}$) aligned with the sequence of the basis. The form of field after transformation will always be two $L_{z}\times L_{z}$ matrix fields differing by $z$-parity symmetry labels, with $S$ counting for symmetric distribution and $A$ for the opposite. Each matrix field will also be attached with a new $4\times 4$ Dirac matrix.
To express matrix quantities $\mathbf{I},\mathbf{G},\mathbf{O},\mathbf{S}$ in Table 1, we introduce the momentum-dependent matrix-form acting functional $\mathcal{F}_{k}$ to represent them over $V$ field that generates projected matrix component like
@@
different by symmetry requirement and an inner sign. In the table the symmetry labels between the transformed fields and the summation kernels are in one-to-one correspondence. For Zeeman field, we have orbital-homogeneous part and orbital-differential parts considering possibly different $g$ factors on basis orbitals. It is worth noting that for ++ kernel, the ‘$A$’ matrix is always obtained from the ‘$S$’ matrix with multiplication $e^{i\pi}\tau_{y}$, while for - kernel, the ‘$A$’ matrix is obtained from the ‘$S$’ matrix with multiplication $e^{-i\pi/2}\tau_{y}$. This interesting phenomenon should trace back to the chiral nature of the Dirac fermions represented by the $\tau_{y}$ matrix (see Appendix A).
This procedure above is general, powerful while easy to understand. Despite the easiness of the transformation, the non-trivial difficult part is to endow physical meaning to the attached fields, both before transformation and after. For instance, the spin-orbital coupling remains its meaning after the transformation, while being block-diagonal in the Dirac fermion representation; the $z$-Zeeman field, as discussed above, is transformed into two matrix form Higgs fields, which stand as the effective mass generators.
@@ V.1 Half quantum mirror Hall effect: a non-magnetic film with mirror symmetry
where $\bar{v}_{i}^{mn,\chi}=\braket{n^{\chi}|\partial_{k_{i}}H^{\chi}|m^{\chi}}$ is the expectation value of the mirror velocity operator evaluated over eigenstates of the mirror-projected Hamiltonian. Clearly, this is just the usual Kubo formula[62] evaluated over the projected Hamiltonian $H_{\chi}$, and thanks to the imposed mirror symmetry, two parts with mirror label $\chi=\pm$ do not communicate with each other and are totally decoupled.
@@
To show the nature of the half quantum mirror Hall effect, we calculate the Hall conductivity of $H_{\chi}$ obtained from the mirror-projected TI film Hamiltonian, and of the split Dirac fermion $H_{\text{surf},\chi}$. The results are shown in Fig. 11, where the half-quantized transverse conductivity nature is shown for each $\chi$ part with inverse signs, indicating quantum spin Hall like physics[12, 82, 43, 44, 45, 83, 84, 85, 86], while the topological origin of the half-quantized mirror Hall conductivity is bound with the metallic gapless Dirac fermions[33]. Their massless low-energy parts distribute mirror-(anti-)symmetrically at both top and bottom surfaces of the TI film as a result from the bulk-boundary correspondence of 3D strong topological insulator[17], corresponding to states with mass $\pm\Theta(-m_{0}(\bm{k}))m_{0}(\bm{k})$ term at $m_{0}(\bm{k})>0$. Here, the symmetry statement is traced back to our basis, which is chosen to distribute along $z$ either mirror symmetrically or anti-symmetrically (see Appendix B). As a complete band, the surface Dirac cone does not end at a finite wavevector, but gradually emerges into the bulk with a regulated non-zero mass term represented by $\Theta(-m_{0}(\bm{k}))m_{0}(\bm{k})$ at $m_{0}(\bm{k})<0$, and it is this non-vanishing high-energy part that ultimately gives rise to the half-quantized Hall conductivity, as discussed in Section III, which finally reads by Eq. (40) as $\sigma_{H}^{\chi}=-\chi e^{2}/2h$, when the Fermi surface satisfies that $m_{0}(k_{F})>0$.
@@
and equals to quantum unit $e^{2}/h$ in the case. The quantity reveals that, though, by opposite Hall conductivity, the charge current by a transverse electrical field vanishes as $\sigma_{H}=\sum_{\chi}\sigma_{H}^{\chi}=0$, the ‘mirror’ current does not, similar to that in quantum spin Hall effect. Nevertheless, a better way of looking at the half quantum mirror Hall effect may start from treating it as an intrinsic ‘spin’ Hall effect in metal, while the effect shows quantization with its transverse ‘spin’ Hall conductivity that shares a topological origin deeply related to the parity anomaly, and replacing ‘spin’ with ‘mirror’ leads to the observation that in different mirror sectors, the mirror current and the charge current will be either parallel or anti-parallel with the same quantized magnitude. Such a way of narration also lies in the lineage of induced dissipationless mirror current and dissipative longitudinal current, as they are both generated by metallic gapless Dirac fermions. To detect the mirror current, non-local electrical transport signals[87, 88, 89] are needed, while to reveal the quantized nature, one needs to perform a series of measurements to fully separate the dissipationless and dissipative currents[33], by changing the sample width and noticing the scale invariance of the Hall conductance.
@@ V.2 Quantum anomalous Hall effect: Chern Insulators
Figure 13: Chern insulator case I: total layer number $L_{z}=19$ with uniform Zeeman field $V_{z}\equiv 0.1\penalty 10000\ $\mathrm{e}\mathrm{V}$\$. (a) Comparison of band structure from TI film model and effective four-band Hamiltonian. (b) Calculated $I_{S}(\bm{k})$ and $I_{A}(\bm{k})$. (c) Calculated Hall conductivity from TI film model and effective four-band Hamiltonian. (d) Hall conductivity for $\chi=\pm$.
Figure 14: Chern insulator case II: total layer number $L_{z}=19$ with symmetric Zeeman field $V_{z}(l_{z})=0.1\penalty 10000\ $\mathrm{e}\mathrm{V}$\$ at top and bottom 22 layers. (a) Comparison of band structure from the lowest four bands of TI film model and effective four-band Hamiltonian. (b) Calculated $I_{S}(\bm{k})$ and $I_{A}(\bm{k})$. (c) Calculated Hall conductivity from TI film model and effective four-band Hamiltonian. (d) Hall conductivity for $\chi=\pm$.
Figure 15: Chern insulator case III: total layer number $L_{z}=19$ with top-22-layer Zeeman field $V_{z}^{t}=0.1\penalty 10000\ $\mathrm{e}\mathrm{V}$\$ and bottom-22-layer field $V_{z}^{b}=0.2\penalty 10000\ $\mathrm{e}\mathrm{V}$\$. (a) Comparison of band structure from the lowest four bands of TI film model and effective four-band Hamiltonian. (b) Calculated $I_{S}(\bm{k})$ and $I_{A}(\bm{k})$. (c) Calculated Hall conductivity from TI film model and effective four-band Hamiltonian. (d) Hall conductivity for $\chi=\pm$.